OpenStudy (anonymous):

I just dont get this at all Maria is debating between two different mortgages for $155,000. She found a 20-year fixed rate loan at 7.35% and 15-year fixed rate loan at the same rate. How much more interest will she pay for the 20-year loan versus the 15-year loan $2,492.45 $40,013.40 $57,737.40 $5,158.30

OpenStudy (dumbcow):

first determine what the monthly payments will be for each loan. use financial calculator or present value equation

OpenStudy (anonymous):

thats what i keep trying to do i just dont understand what im doing wrong.

OpenStudy (dumbcow):

what math class is this for?

OpenStudy (anonymous):

advanced algebra with financial apps

OpenStudy (dumbcow):

ok guess you need to do it by hand here is present value equation which is a geometric sequence i = monthly int rate v = 1/(1+i) 20 yr loan (240 months) \[155,000 = P(v + v^{2}+...v^{240})\] 15 yr loan (180 months) \[155,000 = P(v + v^{2}+...v^{180})\] use formula for sum of geometric series \[S_n = v +v^{2}+...v^{n} = \frac{v(1-v^{n})}{1-v}\] Now you can find the monthly payments for each loan multiply the payment by length of loan to get total payment, subtract loan amount to get total interest paid

OpenStudy (dumbcow):

P is monthly payment

OpenStudy (anonymous):

its b.

OpenStudy (dumbcow):

ok, are you asking or telling me...i haven't worked out the answer yet

OpenStudy (dumbcow):

yes its B do you understand how to figure it out now?

OpenStudy (anonymous):

yes thanks i have another question tho.

OpenStudy (anonymous):

Will is buying a house for $185,000. He is financing $165,000 and obtained a 30-year, fixed-rate mortgage with a 6.725% interest rate. How much are his monthly payments? $506.25 $1,067.45 $1,196.83 $1,656.73

OpenStudy (dumbcow):

same thing as before...except number of months is 360 PV = 165,000 i = 6.725% /12 v = 1/(1+i) \[165,000 = P(v+v^{2}+...v^{360})\] solve for P using the Sum formula

OpenStudy (anonymous):

Correct answer: $1,067.45